# Parafermions with symmetry-protected non-Abelian statistics

### 14 Mar 2024 v1

@@ IV.1 Model with mirror symmetry
We start with a non-interacting model hosting pairs of symmetry-protected MZMs and then switch to the strongly-correlated cases. As depicted in Fig. 4(a), the model consists of two spin-orbit coupled nanowires proximited to two conventional superconductors with $\pi$-phase difference and the Bloch Hamiltonian reads
@@
$\alpha$ $=$ $e^{i[(\varphi_{11\bar{1}}-\varphi_{\bar{1}11})+(\varphi_{111}+\varphi_{\bar{1}1\bar{1}})]/[2(2n+1)]},$ (27)
$\beta$ $=$ $e^{i[(\varphi_{1\bar{1}1}-\varphi_{\bar{1}\bar{1}\bar{1}})+(\varphi_{1\bar{1}\bar{1}}+\varphi_{\bar{1}\bar{1}1})]/[2(2n+1)]}.$ (28)
@@
IV.2 Local mixing under symmetric perturbation
As mentioned in the generic theory, the parity conservation of each sector does not imply the decomposition of the whole system into two distinct subsystems. To show that certain interactions can mix modes in two sectors without ruining non-Abelian statistics, we model the above system with a lattice Hamiltonian (see Fig. 4(d))
@@ IV.2 Local mixing under symmetric perturbation
where $H_{X}=\sum_{i=1}^{N-1}JX_{2i}^{\dagger}X_{2i+1}+J^{*}X_{2i+1}^{\dagger}X_{2i}$ and $\alpha_{i}$, $\beta_{i}$ are parafermionic operators that satisfy $\mathcal{M}\alpha_{i}\mathcal{M}^{-1}=\alpha_{i}$, $\mathcal{M}\beta_{i}\mathcal{M}^{-1}=\omega\beta_{i}$. The third part $H_{p}$ includes certain symmetric perturbation terms $H_{p}=f\alpha_{1}^{\dagger}\alpha_{2}+f^{*}\alpha_{2}^{\dagger}\alpha_{1}+V\alpha_{2}^{\dagger}\alpha_{3}\beta_{2}^{\dagger}\beta_{3}+V^{*}\beta_{3}^{\dagger}\beta_{2}\alpha_{3}^{\dagger}\alpha_{2}$, where $f/J\ll 1$, $V/J\ll 1$, $f/V\sim\mathcal{O}(1)$ and $[H_{p},\mathcal{M}]=0$. In the absence of $H_{p}$, the lattice system is in a dimerized phase and $\alpha_{1,N}$, $\beta_{1,N}$ are exact PZMs. The terms in $H_{p}$ only perturb a single zero mode $\alpha_{1}$ and the new PZM in the presence of $H_{p}$ can be constructed as $\bar{\alpha}_{1}=\alpha_{1}+\sum_{i}m_{i}\hat{O}_{i}^{(1)}+\sum_{j}n_{j}\hat{O}_{j}^{(2)}+\cdots$, where $m_{i}\sim\mathcal{O}(f/J)$ and $n_{j}\sim\mathcal{O}(f^{2}/J^{2})$. By calculating the perturbation series order by order, we find $\tilde{\alpha}_{1}$ cannot be constructed solely by terms with $\alpha_{i}$ , and its dynamics is governed by the full Hamiltonian rather than just $H_{\alpha}$ (see Appendix C). If $\alpha_{i}$ and $\beta_{i}$ are reduced to Majorana operators $\gamma_{i}^{\alpha}$ and $\gamma_{i}^{\beta}$, the exact zero mode can be constructed explicitly
@@
where $a\sim\mathcal{O}(f/J)$, $b\sim\mathcal{O}(f^{2}/J^{2})$ and $c\sim\mathcal{O}(f^{3}/J^{3})$. From this expression, we find the Majorana operators $\beta_{2,3}$ in $H_{\beta}$ are involved in the wavefunction of $\bar{\gamma}_{1}^{\alpha}$, which is an MZM in $\alpha$-sector since $\mathcal{M}\bar{\gamma}_{1}^{\alpha}\mathcal{M}^{-1}=\bar{\gamma}_{1}^{\alpha}$. This confirms that interactions modify the concrete form of the braiding but not affecting the SPNA statistics in general.
@@ V Conclusions and Discussions
This study advances the SPNA statistics from free-fermion topological states to a new realm of strongly correlated topological phases whose classifications are much broader Wen2017Zoo . Our prediction broadly expands the basic notion of non-Abelian statistics in combining with symmetry protection and with or without strong correlations. Many future important issues deserve the further in-depth study. For instance, while in the present study we have focused on the PZMs protected by Abelian unitary symmetries, extending the results to the scenarios with non-Abelian symmetries, which can protect coexisting PZMs more than two, may generate fundamentally new type of SPNA statistics due to the nonlinear features in the fractional quasiparticles, possibly achieving the universal quantum computation. On the other hand, extending SPNA statistics to higher-dimensional non-Abelian topological states CChan2017 ; ZCGu2021 may yield the non-Abelian anyons beyond zero dimension.

### 21 Mar 2024 v2

@@ IV.1 Model with mirror symmetry
We start with a non-interacting model hosting pairs of symmetry-protected MZMs and then turn to the strongly-correlated regime. As depicted in Fig. 4(a), the model consists of two spin-orbit coupled nanowires proximited to two conventional superconductors with $\pi$-phase difference and the Bloch Hamiltonian reads
@@
$\begin{split}\alpha&=&e^{i[(\varphi_{11\bar{1}}-\varphi_{\bar{1}11})+(\varphi_{111}+\varphi_{\bar{1}1\bar{1}})]/[2(2n+1)]},\\ \beta&=&e^{i[(\varphi_{1\bar{1}1}-\varphi_{\bar{1}\bar{1}\bar{1}})+(\varphi_{1\bar{1}\bar{1}}+\varphi_{\bar{1}\bar{1}1})]/[2(2n+1)]}.\end{split}$ (29)
@@
IV.2 Effect of nonlinear interactions
As mentioned in the generic theory, the parity conservation of each sector does not imply the decomposition of the whole system into two decoupled subsystems. To show that certain interactions can mix modes in two sectors without ruining non-Abelian statistics, we model the above system with a lattice Hamiltonian (see Fig. 4(d))
@@ IV.2 Effect of nonlinear interactions
where $H_{X}=\sum_{i=1}^{N-1}JX_{2i}^{\dagger}X_{2i+1}+J^{*}X_{2i+1}^{\dagger}X_{2i}$ and $\alpha_{i}$, $\beta_{i}$ are parafermionic operators that satisfy $\mathcal{M}\alpha_{i}\mathcal{M}^{-1}=\alpha_{i}$, $\mathcal{M}\beta_{i}\mathcal{M}^{-1}=\omega\beta_{i}$. The third part $H_{p}$ includes certain symmetric perturbation terms $H_{p}=f\alpha_{1}^{\dagger}\alpha_{2}+f^{*}\alpha_{2}^{\dagger}\alpha_{1}+V\alpha_{2}^{\dagger}\alpha_{3}\beta_{2}^{\dagger}\beta_{3}+V^{*}\beta_{3}^{\dagger}\beta_{2}\alpha_{3}^{\dagger}\alpha_{2}$, where $f/J\ll 1$, $V/J\ll 1$, $f/V\sim\mathcal{O}(1)$ and $[H_{p},\mathcal{M}]=0$. In the absence of $H_{p}$, the lattice system is in a dimerized phase, with $\alpha_{1,2N}$, $\beta_{1,2N}$ being exact parafermion zero edge modes, which characterizes the results in the above nanowire model. The perturbation terms in $H_{p}$ are only introduced to the zero mode $\alpha_{1}$ and the new PZM in the presence of $H_{p}$ can be constructed as $\bar{\alpha}_{1}=\alpha_{1}+\sum_{i}m_{i}\hat{O}_{i}^{(1)}+\sum_{j}n_{j}\hat{O}_{j}^{(2)}+\cdots$, where $m_{i}\sim\mathcal{O}(f/J)$ and $n_{j}\sim\mathcal{O}(f^{2}/J^{2})$. By calculating the perturbation series order by order, we find $\tilde{\alpha}_{1}$ cannot be constructed solely by terms with $\alpha_{i}$ , and its dynamics is governed by the full Hamiltonian rather than just $H_{\alpha}$ (see Appendix C). For an intuitive result, if $\alpha_{i}$ and $\beta_{i}$ are reduced to Majorana operators $\gamma_{i}^{\alpha}$ and $\gamma_{i}^{\beta}$, the exact zero mode can be constructed explicitly
@@
where $a\sim\mathcal{O}(f/J)$, $b\sim\mathcal{O}(f^{2}/J^{2})$ and $c\sim\mathcal{O}(f^{3}/J^{3})$. From this expression, we see that the Majorana modes $\beta_{2,3}$ in $H_{\beta}$ are involved in the wavefunction of $\bar{\gamma}_{1}^{\alpha}$, which is an MZM in $\alpha$-sector since $\mathcal{M}\bar{\gamma}_{1}^{\alpha}\mathcal{M}^{-1}=\bar{\gamma}_{1}^{\alpha}$. This result confirms that interactions modify the concrete form of the zero modes and the effective diagonal bases of the braiding, but not affecting the SPNA statistics since the generic symmetry protection mechanism is preserved.
@@ V Conclusions and Discussions
This study advances the SPNA statistics from free-fermion topological states to a new realm of strongly correlated topological phases whose classifications are much broader Wen2017Zoo . Our prediction broadly expands the basic notion of non-Abelian statistics in combining with symmetry protection and strong correlations. A lot of future important issues deserve further in-depth study. For instance, while in the present study we have focused on the PZMs protected by Abelian unitary symmetries, extending the results to the scenarios with non-Abelian symmetries, which can protect coexisting PZMs more than two, may generate fundamentally new type of SPNA statistics due to the nonlinear features in the fractional quasiparticles, possibly achieving the universal quantum computation. On the other hand, extending SPNA statistics to higher-dimensional non-Abelian topological states CChan2017 ; ZCGu2021 may yield the symmetry-protected non-Abelian anyons in beyond one and two dimensions.